# R&D Networks under Heterogeneous Firm Productivities

### v1

@@ 2.1 Equilibrium Characterization
These last two stages can be interpreted as the solution to a game with an exogenously given R&D network (a similar structure is analyzed in konig2019r, under the assumptions of homogeneous productivity and multiple markets). Finally, we solve for the equilibrium network structure invoking the concept of pairwise stability (jackson1996strategic), characterizing the endogenous formation of R&D alliances for heterogeneous firms.
@@ 2.1.1 Market Competition
$\begin{aligned} {q}^{*}_{i}\left(\mathbf{e},\mathbf{G}\right)&=\frac{1}{n+1}\left[\alpha-n\cdot c_{i}\left(\mathbf{e},\mathbf{G}\right)+\sum_{k\neq i}c_{k}\left(\mathbf{e},\mathbf{G}\right)\right],\\ {\pi}^{*}_{i}\left(\mathbf{e},\mathbf{G}\right)&={q}^{*}_{i}\left(\mathbf{e},\mathbf{G}\right)^{2}-\phi\cdot e_{i}^{2}.\end{aligned}$ (4)
@@
where $\mathcal{N}_{-i}(\mathbf{G})\coloneqq\left\{j\in\mathcal{N}\;|\;G_{ij}=0\right\}$ denotes the set of firms that are not neighbors of firm $i$ in the network $\mathbf{G}$. Equation (5) shows that a firm’s optimal production level increases with its own R&D effort and decreases with the R&D efforts of its non-neighbors. The effect of R&D efforts by firm $i$’s neighbors on its production is positive given that the degree of each firm $j\in\mathcal{N}$ always satisfies $0\leq d_{j}(\mathbf{G})\leq n-1<n$.
@@ 2.1.2 Equilibrium R&D Efforts
$\left.\frac{\partial{\pi}^{*}_{i}\left(\mathbf{e},\mathbf{G}\right)}{\partial e_{i}}\right|_{{e}^{*}_{i}}=\quad\frac{2{q}^{*}_{i}\left(\mathbf{e},\mathbf{G}\right)}{n+1}\left(-n\left.\frac{\partial c_{i}\left(\mathbf{e},\mathbf{G}\right)}{\partial e_{i}}\right|_{{e_{i}}^{*}}+\sum_{k\neq i}\left.\frac{\partial c_{k}\left(\mathbf{e},\mathbf{G}\right)}{\partial e_{i}}\right|_{{e_{i}}^{*}}\right)-2\phi\;{e}^{*}_{i}=0,$
where ${e}^{*}_{i}\left(\mathbf{e}_{-i},\mathbf{G}\right)\coloneqq\arg\max_{e_{i}}\{{\pi}^{*}_{i}\}$ is firm $i$’s best-response effort level, and $\mathbf{e}_{-i}\coloneqq\{e_{j}\}_{j\neq i}$ is the effort of all other firms. Let $\eta_{i}\left(\mathbf{G}\right)\coloneqq[n-d_{i}\left(\mathbf{G}\right)]/(n+1)$ represent a strictly positive sparsity coefficient that decreases with the firm’s degree $d_{i}(\mathbf{G})$. From equation (3), we know that $\partial c_{i}(\mathbf{e},\mathbf{G})/\partial e_{i}=-\theta_{i}$ and $\partial c_{i}(\mathbf{e},\mathbf{G})/\partial e_{j}=-\theta_{j}\,G_{ij}$, such that we can rewrite the FOC as:
${e}^{*}_{i}\left(\mathbf{G},\mathbf{e}_{-i}\right)=\underbrace{\frac{\theta_{i}\;\eta_{i}(\mathbf{G})}{\phi-\theta_{i}^{2}\;\eta_{i}^{2}(\mathbf{G})}}_{\text{\scriptsize(i) sparsity}}\cdot\Biggl[\underbrace{\frac{\alpha-\bar{c}}{n+1}}_{\text{\scriptsize(ii) baseline}}+\underbrace{\sum_{j\in\mathcal{N}_{i}(\mathbf{G})}\eta_{j}(\mathbf{G})\;\theta_{j}\;e_{j}}_{\text{\scriptsize(iii) neighbor effect}}-\underbrace{\sum_{k\in\mathcal{N}_{-i}(\mathbf{G})}\big[1-\eta_{k}(\mathbf{G})\big]\;\theta_{k}\;e_{k}}_{\text{\scriptsize(iv) non-neighbor effect}}\Biggr],$ (6)
The best-response effort of firms can be decomposed into four key factors. The first is the proportionality constant that is increasing in the firm’s productivity $\theta_{i}$, and its sparsity coefficient $\eta_{i}(\mathbf{G})$, and therefore, decreasing in the degree $d_{i}(\mathbf{G})$. The second is a baseline effort level that depends on market parameters $\alpha$, $\bar{c}$, and $n$. The structures of the neighbor effect (iii) and non-neighbor effects (iv) mirror that in equation (5), indicating that a firm’s optimal R&D effort is increasing in neighbors’ efforts and productivities, and decreasing in neighbors’ degrees as well as in non-neighbors’ efforts, productivities and degrees.
@@
$\mathbf{A}(\mathbf{G})\;\mathbf{e}=(\alpha-\bar{c})\;\mathbf{1}_{n},\quad\text{ where }\quad A_{ij}(\mathbf{G})=\begin{cases}\frac{(n+1)^{2}\;\phi}{\theta_{i}\;[n-d_{i}(\mathbf{G})]}-\theta_{i}\;[n-d_{i}(\mathbf{G})]&\text{for }i=j,\\ \left[1+d_{j}(\mathbf{G})\right]\;\theta_{j}-(n+1)\;G_{ij}\;\theta_{j}&\text{for }j\neq i.\end{cases}$ (7)
Proposition 1.
There exists a unique equilibrium effort profile ${\mathbf{e}}^{*}(\mathbf{G})=\left[{e}^{*}_{1}(\mathbf{G}),\cdots,{e}^{*}_{n}(\mathbf{G})\right]^{\top}$ with ${e}^{*}_{i}(\mathbf{G})>0$ for all $i\in\mathcal{N}$ that solves equation (7) if the cost parameter $\phi$ satisfies:
$\phi>\underline{\phi}\coloneqq\frac{n^{3}-n^{2}-n+2}{n^{2}+2n+1}.$ (8)
Proposition 1 establishes that a unique and interior equilibrium exists for any network $\mathbf{G}\in\mathcal{G}^{n}$, provided the cost coefficient $\phi$ exceeds a threshold $\underline{\phi}$ that grows linearly with $n$. The proof of Proposition 1 relies on the Gershgorin Disk Theorem (gershgorin1931uber). Specifically, we show that the matrix $\mathbf{A}$ is strictly diagonally dominant, which implies that it is non-singular and hence the FOC system admits a unique solution.
While existence and uniqueness of an equilibrium effort profile are guaranteed by Proposition 1, the equilibrium effort levels cannot be generally expressed in closed form, as they involve an intractable inverse of the non-symmetric matrix $\mathbf{A}(\mathbf{G})$. Nevertheless, we can derive equilibrium profits using equilibrium efforts as:
${\pi}^{*}_{i}\left(\mathbf{G}\right)=\left[\frac{\phi}{\theta_{i}^{2}\;\eta_{i}^{2}(\mathbf{G})}\;-1\right]\;\phi\;{{e}^{*}_{i}}^{2}\left(\mathbf{G}\right).$ (9)
@@
For two firms $i,j$ that have a symmetric position, they hold identical links in the network $\mathbf{G}$. This directly implies their degrees are also the same, $d_{i}(\mathbf{G})=d_{j}(\mathbf{G})$, and so are their sparsity coefficients $\eta_{i}(\mathbf{G})=[n-d_{i}(\mathbf{G})]/(n+1)=[n-d_{j}(\mathbf{G})]/(n+1)=\eta_{j}(\mathbf{G})$.
Proposition 2.
@@
This can be compared with the result in hsieh2024endogenous, showing ${e}^{*}_{i}=\theta{q}^{*}_{i}/(2\phi)$, with the relevant changes in notation. The differences in equilibrium effort and profit levels in these two setups arise from, first, differences in sources of heterogeneity (in baseline costs in their model vs. R&D productivities in our model) and also the solution methods implemented for equilibrium construction. Equation (11) implies that a firm’s equilibrium size is positively related to its equilibrium R&D effort, consistent with empirical findings in cohen1996firm. Additionally, we see from the expression that the optimal output level is negatively related to its productivity $\theta_{i}$ and its sparsity coefficient $\eta_{i}(\mathbf{G})$. Equations (10) and (11) imply that for $\theta_{i}>\theta_{j}$:
@@ 2.1.3 R&D Network Formation
Pairwise stability requires that no firm has an incentive to sever an existing link, and that no two firms both benefit (with at least one strictly gaining) from forming a new one (jackson1996strategic). In our case, the formal statement translates to the following: a network $\mathbf{G}$ is pairwise stable if and only if for all firms $i,j\in\mathcal{N}$,
  • Existing Alliances: If ${G}_{ij}=1$, then neither firm $i$ nor firm $j$ would benefit from dissolving the alliance, i.e., $\pi_{i}(\mathbf{G})\geq\pi_{i}(\mathbf{G}_{-ij})$ and $\pi_{j}(\mathbf{G})\geq\pi_{j}(\mathbf{G}_{-ij})$.
@@
Let the complete network $\mathbf{G}^{C}$ be the one in which all firms are linked to every other firm ($\forall i\neq j\in\mathcal{N},G^{C}_{ij}=1$). Under homogeneity, goyal2001r shows that the complete network is pairwise stable. By generalizing to heterogeneous R&D productivities, we present our next result.
Proposition 3.
@@
There exists $\theta^{\star}\in(0,1)$ such that if there is a firm $j$ with $\theta_{j}=\theta^{\star}$, then the complete network $\mathbf{G}^{C}$ is not pairwise stable.
Proposition 3 shows that the complete network, which is stable under homogeneous productivities, need not remain stable when productivities are heterogeneous. When a sufficiently large productivity gap exists between the most productive firm (say $i$) and any competitor $j$, a link between them cannot be sustained, leading to instability of the complete network. The proof of Proposition 3 proceeds by comparing the equilibrium profit of the most productive firm $i$ (defined as having $\theta_{i}=1$) in the complete network, ${\pi}^{*}_{i}(\mathbf{G}^{C}_{+ij})$, vs. in the network that severs the link between $i$ and $j$, ${\pi}^{*}_{i}(\mathbf{G}^{C}_{-ij})$. We show that removing the link is profitable for the most productive firm as $\theta_{j}$ approaches 0. Conversely, when $\theta_{j}$ approaches 1, such a deviation is not profitable and the complete network is pairwise stable.
Finally, through simulations, we show that the relative profit change, $\Delta^{ij}{\pi}^{*}_{i}(\mathbf{G}^{C})/{\pi}^{*}_{i}(\mathbf{G}^{C}_{-ij})$, is monotonically increasing in $\theta_{j}$ (R&D productivity of potential partner): a low-type firm $j$ always benefits from creating the connection to the most productive firm $i$, whereas firm $i$ has a threshold level $\theta^{*}$ such that it would deviate (increase its profits) by removing the link when $\theta_{j}<\theta^{*}$. This suggests the stronger result that, for any given network $\mathbf{G}$, a link $G_{ij}$ between the most productive firm $i$ and a competitor $j$ cannot be sustained if there is a sufficiently large productivity gap $\theta_{j}<{\theta}^{*}$.
@@
To obtain these insights, in Figure 1, we simulate random networks of $n=20$ firms using an Erdos-Renyi generation (erdHos1960evolution), with probability of connection $\ell\in[0,1]$ (where $\ell=0$ corresponds to the empty network and $\ell=1$ to the complete network). We obtain productivity distributions $(\theta_{1},\ldots,\theta_{n})$ as random samples from Beta distributions with varying parameters. Finally, we fix the productivities of two firms $i$ and $j$, having $\theta_{i}=1$ and allow $\theta_{j}$ to vary between 0 and $1$, and show how adding/severing the link between such $i$ and $j$ affects their profits, while changing the density of the network and the productivity distribution for (other) firms. Figure 1 plots the productivity distributions in the upper panel that are used to calculate profit changes $\Delta^{ij}{\pi}^{*}(\mathbf{G})/{\pi}^{*}(\mathbf{G}_{-ij})$ in the main (lower) panel, also varying the level of network connectivity $\ell$ within each case.
@@
Figure 1: Percentage changes in profits of a high- and low-productivity firm after forming an R&D collaboration link between them in random networks with varying connection density and firm productivity distributions
@@
Note: Networks $\mathbf{G}$ with $n=20$ firms generated as random binary adjacency matrices according to an Erdos-Renyi scheme with connectivity parameter $\ell\in\{0,\;0.25,\;0.75,\;1\}$, represented by different colors. Productivity distributions $(\theta_{1},\ldots,\theta_{n})$ drawn as random samples from a Beta(0.5,0.5)(0.5,0.5) distribution (left); Beta(1,1)(1,1) (middle); and Beta(2,2)(2,2) (right), with corresponding density plotted in the panels above. The percentage change in profit of firm $i$ having $\theta_{i}=1$ (before and after adding a link to firm $j$) is drawn using a solid line, and for firm $j$ using a dashed line. Percentage change in profits for firm $i$ crosses 0 at a threshold ${\theta}^{*}\in(0,1)$ for each productivity distribution and connectivity level cases shown.
Next, we provide further simulation-based findings in Figure 2 showing that for an arbitrary firm $i$ with productivity $\theta_{i}$, there exists a threshold $0<{\theta}^{*}_{i}<\theta_{i}$, such that firm $i$ would not form a collaboration with a firm $j$ having $\theta_{j}<{\theta}^{*}_{i}$; i.e., if firm $j$ is sufficiently less productive than firm $i$. This generalizes the insights in Proposition 3 and Figure 1, shown only for the most productive firm, to any pair of randomly selected firms with arbitrary productivities.
Figure 2: Extension of Figure 1 to a general setting with two firms $i$ and $j$ having arbitrary productivities $\theta_{i}>\theta_{j}$
Note: Simulation setting is identical to that of Figure 1, except the productivity of the higher-productive firm $i$ is chosen from $\theta_{i}\in\{0.25,0.5,0.75\}$. Figures are limited to the range $0<\theta_{j}<\theta_{i}$ of each configuration to highlight the crossing at 0 for the higher-productive firm $i$.

### v2

@@ 2.1 Equilibrium Characterization
These last two stages can be interpreted as the solution to a game with an exogenously given R&D network (a similar structure is analyzed in König et al., 2019, under the assumptions of homogeneous productivity and multiple markets). Finally, we solve for the equilibrium network structure invoking the concept of pairwise stability (Jackson and Wolinsky, 1996), characterizing the endogenous formation of R&D alliances for heterogeneous firms.
@@ 2.1.1 Market Competition
$\begin{aligned} {q}^{*}_{i}\left(\mathbf{e},\mathbf{G}\right)&=\frac{1}{n+1}\left[\alpha-n\cdot c_{i}\left(\mathbf{e},\mathbf{G}\right)+\sum_{j\neq i}c_{j}\left(\mathbf{e},\mathbf{G}\right)\right],\\ {\pi}^{*}_{i}\left(\mathbf{e},\mathbf{G}\right)&={q}^{*}_{i}\left(\mathbf{e},\mathbf{G}\right)^{2}-\phi\cdot e_{i}^{2}.\end{aligned}$ (4)
@@
where $\mathcal{N}_{-i}(\mathbf{G})\coloneqq\left\{j\in\mathcal{N}\mid G_{ij}=0,j\neq i\right\}$ denotes the set of firms that are not neighbors of firm $i$ in the network $\mathbf{G}$. Equation (5) shows that a firm’s optimal production level increases with its own R&D effort and decreases with the R&D efforts of its non-neighbors. The effect of R&D efforts by firm $i$’s neighbors on its production is positive given that the degree of each firm $j\in\mathcal{N}$ always satisfies $0\leq d_{j}(\mathbf{G})\leq n-1<n$, while the effect from non-neighbours is therefore always negative.
@@ 2.1.2 Equilibrium R&D Efforts
$\left.\frac{\partial{\pi}^{*}_{i}\left(\mathbf{e},\mathbf{G}\right)}{\partial e_{i}}\right|_{{e}^{*}_{i}}=\quad\frac{2{q}^{*}_{i}(\mathbf{e},\mathbf{G})}{n+1}\left(-n\left.\frac{\partial c_{i}(\mathbf{e},\mathbf{G})}{\partial e_{i}}\right|_{{e_{i}}^{*}}+\sum_{j\neq i}\left.\frac{\partial c_{j}(\mathbf{e},\mathbf{G})}{\partial e_{i}}\right|_{{e_{i}}^{*}}\right)-2\phi{e}^{*}_{i}=0,$
where ${e}^{*}_{i}\left(\mathbf{e}_{-i},\mathbf{G}\right)\coloneqq\arg\max_{e_{i}}\{{\pi}^{*}_{i}\}$ is firm $i$’s best-response effort level, and $\mathbf{e}_{-i}\coloneqq\{e_{j}\}_{j\neq i}$ is the effort of all other firms. Let $\eta_{i}\left(\mathbf{G}\right)\coloneqq[n-d_{i}\left(\mathbf{G}\right)]/(n+1)$ represent a strictly positive sparsity coefficient that decreases with the firm’s degree $d_{i}(\mathbf{G})$. From equation (3), we know that $\partial c_{i}(\mathbf{e},\mathbf{G})/\partial e_{i}=-\theta_{i}$ and $\partial c_{j}(\mathbf{e},\mathbf{G})/\partial e_{i}=-\theta_{i}G_{ij}$, such that we can solve the FOC in order to obtain:
${e}^{*}_{i}\left(\mathbf{G},\mathbf{e}_{-i}\right)=\underbrace{\frac{\theta_{i}\eta_{i}(\mathbf{G})}{\phi-\theta_{i}^{2}\eta_{i}^{2}(\mathbf{G})}}_{\text{\scriptsize(i) sparsity}}\Biggl\{\underbrace{\frac{\alpha-\bar{c}}{n+1}}_{\text{\scriptsize(ii) baseline}}+\underbrace{\sum_{j\in\mathcal{N}_{i}(\mathbf{G})}\eta_{j}(\mathbf{G})\theta_{j}e_{j}}_{\text{\scriptsize(iii) neighbor effect}}-\underbrace{\sum_{k\in\mathcal{N}_{-i}(\mathbf{G})}\big[1-\eta_{k}(\mathbf{G})\big]\theta_{k}e_{k}}_{\text{\scriptsize(iv) non-neighbor effect}}\Biggr\},$ (6)
The best-response effort of firms can be decomposed into four key factors. The first is a proportionality constant that is increasing in the firm’s productivity $\theta_{i}$ and its sparsity coefficient $\eta_{i}(\mathbf{G})$, and therefore decreasing in its degree $d_{i}(\mathbf{G})$. The second is a baseline effort level that depends on market parameters $\alpha$, $\bar{c}$, and $n$. The structures of the neighbor effect (iii) and non-neighbor effects (iv) mirror that in equation (5), indicating that a firm’s optimal R&D effort is increasing in neighbors’ efforts and productivities, and decreasing in non-neighbors’ efforts and productivities as well as the degree of all competitor firms.
@@
$\mathbf{A}(\mathbf{G})\mathbf{e}=(\alpha-\bar{c})\mathbf{1}_{n},\quad\text{where }A_{ij}(\mathbf{G})=\begin{cases}\frac{(n+1)^{2}\phi}{\theta_{i}[n-d_{i}(\mathbf{G})]}-\theta_{i}[n-d_{i}(\mathbf{G})]&\text{for }i=j,\\ \left[1+d_{j}(\mathbf{G})\right]\theta_{j}-(n+1)G_{ij}\theta_{j}&\text{for }j\neq i.\end{cases}$ (7)
Proposition 1 (Cost parameter bound).
There exists a unique optimal effort profile ${\mathbf{e}}^{*}(\mathbf{G})\coloneqq\{{e}^{*}_{j}(\mathbf{G})\}_{j=1}^{n}$ with ${e}^{*}_{i}(\mathbf{G})>0$ for all $i\in\mathcal{N}$ that solves equation (7) if the cost parameter $\phi$ satisfies:
$\phi>\underline{\phi}\coloneqq\frac{n^{3}-n^{2}+n}{(n+1)^{2}}.$ (8)
Proposition 1 establishes that a unique and interior equilibrium exists for any network $\mathbf{G}\in\mathcal{G}^{n}$, provided the cost coefficient $\phi$ exceeds a threshold $\underline{\phi}$ that grows linearly with $n$. The proof of Proposition 1 relies on the Levy-Desplanques theorem, itself a consequence of Gershgorin’s disc theorem (Gershgorin, 1931). Specifically, we provide a bound for $\phi$ that guarantees the matrix $\mathbf{A}(\mathbf{G})$ is strictly diagonally dominant regardless of the network $\mathbf{G}$, which implies that it is non-singular and hence the FOC system admits a unique solution.
While existence and uniqueness of an equilibrium effort profile are guaranteed by Proposition 1, equilibrium effort levels cannot be generally expressed in closed form, as they involve an intractable inverse of the non-symmetric matrix $\mathbf{A}(\mathbf{G})$. Nevertheless, we can derive equilibrium profits using equilibrium efforts as:
${\pi}^{*}_{i}\left(\mathbf{G}\right)=\left[\frac{\phi}{\theta_{i}^{2}\eta_{i}^{2}(\mathbf{G})}-1\right]\phi\cdot{{e}^{*}_{i}}^{2}\left(\mathbf{G}\right).$ (9)
@@
For two firms $i$ and $j$ that have a symmetric position, they hold identical links in the network $\mathbf{G}$. This directly implies their degrees are also the same, $d_{i}(\mathbf{G})=d_{j}(\mathbf{G})$, and so are their sparsity coefficients $\eta_{i}(\mathbf{G})=[n-d_{i}(\mathbf{G})]/(n+1)=[n-d_{j}(\mathbf{G})]/(n+1)=\eta_{j}(\mathbf{G})$.
Proposition 2 (Effort and profit ratios under symmetric positions).
@@
This can be compared with the result in Hsieh et al. (2025), showing ${e}^{*}_{i}=\theta{q}^{*}_{i}/(2\phi)$ (with the relevant changes in notation). The differences in equilibrium effort and profit levels in these two setups arise from, first, differences in sources of heterogeneity (in baseline costs in their model vs. R&D productivities in ours) and, second, the solution methods implemented for equilibrium construction (simultaneous in their model vs. sequential action in ours). Equation (11) implies that a firm’s equilibrium size is positively related to its equilibrium R&D effort, consistent with empirical findings in Cohen and Klepper (1996). Additionally, we see from the expression that the optimal output level is negatively related to its productivity $\theta_{i}$ and its sparsity coefficient $\eta_{i}(\mathbf{G})$. Equations (10) and (11) imply that for comparing one high- to one low-productive firm with $\theta_{i}>\theta_{j}$:
@@ 2.1.3 R&D Network Formation
Pairwise stability requires that no firm has an incentive to sever an existing link, and that no two firms both benefit (with at least one strictly gaining) from forming a new one (Jackson and Wolinsky, 1996). In our case, the formal statement translates to the following: a network $\mathbf{G}$ is pairwise stable if and only if for all firms $i,j\in\mathcal{N}$,
  • Existing Alliances: If ${G}_{ij}=1$, then neither firm $i$ nor firm $j$ would benefit from dissolving the alliance, i.e., ${\pi}^{*}_{i}(\mathbf{G})\geq{\pi}^{*}_{i}(\mathbf{G}_{-ij})$ and ${\pi}^{*}_{j}(\mathbf{G})\geq{\pi}^{*}_{j}(\mathbf{G}_{-ij})$.
@@
Let the complete network $\mathbf{G}^{C}$ be the one in which all firms are linked to every other firm ($\forall i,j\in\mathcal{N}$ with $i\neq j$, $G^{C}_{ij}=1$). Under homogeneity, Goyal and Moraga-González (2001) show that the complete network is pairwise stable. By generalizing to heterogeneous R&D productivities, we present our next result.
Proposition 3 (Thresholds for complete network stability).
@@
For any productivity distribution $\{\theta_{i}\}_{i\in\mathcal{N}}\in(0,1]^{n}$ with $\max_{i\in\mathcal{N}}\theta_{i}=1$, there exist $\theta^{*},\theta^{**}\in(0,1)$ with $\theta^{*}\leq\theta^{**}$ such that the complete network $\mathbf{G}^{C}$ satisfies:
  1. (i)
    If $\exists j\in\mathcal{N}$ for which $\theta_{j}<\theta^{*}$, then $\mathbf{G}^{C}$ is not pairwise stable.
  2. (ii)
    If $\theta_{j}\geq\theta^{**}$ for all $j\in\mathcal{N}$, then $\mathbf{G}^{C}$ is pairwise stable.
Proposition 3 shows that when all firms are sufficiently similar in productivity, the complete network remains stable, as in the homogeneous benchmark. However, when productivity differences are large, highly productive firms no longer find it profitable to maintain links with much less productive partners. In this case, the complete network becomes unstable because the most productive firms prefer to sever some of their connections. The key intuition is simple. Linking to a less productive firm yields little benefit for a highly productive firm but still requires sharing the gains from its own R&D effort. As the productivity gap widens, this imbalance grows, eventually making some links unattractive for the most productive firms. Stability therefore breaks down once the productivity of any firm falls below a threshold.
2.2 Simulation preview: Link sustainability for random networks
To generalize insights on pairwise stability beyond the complete network setting in Proposition 3, we provide simulation-based evidence in Figure 1 of link sustainability in random networks. In these simulations, we fix an arbitrary firm $i$ with productivity $\theta_{i}$ and consider a firm $j\neq i$ with a productivity $\theta_{j}$, identifying pair-specific threshold ${\theta}^{*}_{ij}$ values such that $i$ would not form a collaboration with $j$ if $\theta_{j}<{\theta}^{*}_{ij}<\theta_{i}$; i.e., if firm $j$ is sufficiently less productive than firm $i$. These pair-specific threshold values now depend on both the productivity distribution ($\beta$-distribution panels above) and the random network structure $\mathbf{G}$ with varying connection densities shown in different colors.
@@ 2.2 Simulation preview: Link sustainability for random networks
In Figure 1, we simulate random networks of $n=20$ firms using an Erdős-Rényi generation (Erdős and Rényi, 1960), with probability of connection $\ell\in[0,1]$ (where $\ell=0$ corresponds to the empty network and $\ell=1$ to the complete network). We obtain productivity distributions $(\theta_{1},\ldots,\theta_{n})$ as random samples from Beta distributions with varying parameters. Finally, we fix the productivities of two firms $i$ and $j$, having $\theta_{i}\in\{0.25,0.5,0.75\}$ (in three different rows) and allow $\theta_{j}$ to vary between 0 and $\theta_{i}$ ($x$-axes values), and show how severing the link between such $i$ (solid line) and $j$ (dashed line) affects their profits, while changing the density of the network and the productivity distribution for other firms. Figure 1 plots the productivity distributions in the upper panel that are used to calculate profit changes $\Delta^{ij}{\pi}^{*}(\mathbf{G})/{\pi}^{*}(\mathbf{G}_{-ij})$ in the main (lower) panel, also varying the level of network connectivity $\ell$ within each case.
@@
Figure 1: Percentage changes in profits of two arbitrary firms after forming an R&D collaboration link between them in random networks with varying connection density and firm productivity distributions
@@
Note: Networks $\mathbf{G}$ with $n=20$ firms generated as random binary adjacency matrices according to an Erdos-Renyi scheme with connectivity parameter $\ell\in\{0,0.25,0.75,1\}$, represented by different colors. Productivity distributions $(\theta_{1},\ldots,\theta_{n})$ drawn as random samples from a Beta(0.5,0.5)(0.5,0.5) distribution (left); Beta(1,1)(1,1) (middle); and Beta(2,2)(2,2) (right), with corresponding density plotted in the panels above. The percentage change in profit of firm $i$ (before and after adding a link to firm $j$) is drawn using a solid line, and for firm $j$ using a dashed line. Productivity of the higher-productive firm $i$ is chosen from $\theta_{i}\in\{0.25,0.5,0.75\}$. Percentage change in profits for firm $i$ crosses 0 at a threshold ${\theta}^{*}_{ij}\in(0,1)$ for each productivity distribution and connectivity level cases shown. Figures are limited to the range $0<\theta_{j}<\theta_{i}$ of each configuration to highlight the crossing at 0 for the higher-productive firm $i$.